OpenStudy (anonymous):

solve by elimination is anyone good at explaining this? 2x-y+z=-2 x+3y-z=10 x+2z = -8

OpenStudy (turingtest):

first lets number your equations: 2x-y+z=-2 (1) x+3y-z=10 (2) x+2z = -8 (3) we want to get rid of at least one variable. Now notice that (1) has a "+z" and (2) has a "-z", so adding (1) to (2) gives 3x+2y=8 (4) and multiplying (2) by 2, then adding it to (3) gives 3x+6y=12 (5) subtracting (5)-(4) will get rid of the x's: 4y=4 y=1 lets plug that value into (1) and (2). Our modified equations are now 2x-1+z=-2 ->2x+z=-1 (6) x+3-z=10 -> x-z=7 (7) get rid of the z again by adding (6) to (7): 3x=6 x=2 now we can solve any equation for z and plug in the values for x=2 and y=1 to find the value of z. Solving (2) for z gives z=x+3y-10=2+3-10=-5 so x=2 y=1 z=-5 I know that was long and tedious, but try to follow my explanation closely. PS: If (hopefully "when") you take Linear Algebra you will learn far easier and more efficient methods for solving these systems, so hold out!